let f(x)=x^2 -3 and g(x)=2x+3 Evaluate.

f(-2)

Let f(x)=
Answer #1

Soli- f(x) = 2 2-3 g(x)= 2273 Then fl-2) = (-83 4-3 i f(-21 f (g (x)) = 2²8 a=23 2 2 (21+3)-3 7 42+199+93 . f [90] = 47712146